# Are you doing better than random guessing? A call for using negative controls when evaluating causal discovery algorithms

### v1

@@ 7 Discussion
In conclusion, we believe the results and examples provided here showcase that we need to acknowledge that causal discovery is not just another machine learning problem. Estimating a high-dimensional object such as a graph is difficult, and evaluating how well one did is equally challenging. A graph is more than the sum of its edges, and if we do not take into account the most fundamental property of the graphs we simulate for evaluation – its density – we are not producing useful results that will be likely to generalize to new data with other densities. We believe that the use of negative controls will be a useful next step in the direction of more transparent and interpretable evaluations. We of course all hope to do better than random guessing, so let us make it easy to see when we do - and when we do not.
Appendix A: Computations for Table YY
@@
Below, we let $q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$ be the $k$th quantile from $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. We also note, and use repeatedly below, that $TP\operatorname{|}(m_{\text{max}},m_{\text{true}},m_{\text{est}})\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, and hence $\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}$.
@@
We write recall as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
recall $=\frac{TP}{TP+FN}$
$=\frac{TP}{TP+m_{\text{true}}-TP}$
$=\frac{TP}{m_{\text{true}}}$
@@
$\text{E}(\text{recall}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}))$ $=\frac{1}{m_{\text{true}}}\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
@@
$\frac{1}{m_{\text{true}}}q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$
@@
We write the F1 score as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
F1 $=\frac{2\cdot TP}{2\cdot TP+FP+FN}$
$=\frac{2\cdot TP}{2\cdot TP+m_{\text{est}}-TP+m_{\text{true}}-TP}$
$=\frac{2\cdot TP}{m_{\text{est}}+m_{\text{true}}}$
@@
$\frac{2\cdot q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{est}}+m_{\text{true}}}$
@@
We write the negative predictive value (NPV) as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
NPV $=\frac{TN}{TN+FN}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP+FN}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}}$
@@
$\text{E}(\text{NPV}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}))$ $=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})}{m_{\text{max}}-m_{\text{est}}}$
@@
$\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{max}}-m_{\text{est}}}$
@@
We write specificity as a function of $TP,m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
specificity $=\frac{TN}{TN+FP}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP+FP}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{true}}}$
@@
$\text{E}(\text{specificity}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}))$ $=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\text{E}(TP\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})}{m_{\text{max}}-m_{\text{true}}}$
@@
$\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{max}}-m_{\text{true}}}$

### v2

@@ 7 Discussion
In conclusion, we believe the results and examples provided here showcase that we need to acknowledge that causal discovery is not just another machine learning problem. Estimating a high-dimensional object such as a graph is difficult, and evaluating how well one did is equally challenging. If we do not take into account the most fundamental property of the graphs we simulate for evaluation — their densities — we are not producing useful results that will be likely to generalize to new data with other graph densities. We believe that the use of negative controls will be a useful next step in the direction of more transparent and interpretable evaluations. We of course all hope to do better than random guessing, so let us make it easy to see when we do - and when we do not.
Acknowledgements.
The author thanks Vanessa Didelez for her insightful feedback on this work.
Appendix A Random Urn Experiment Motivation for Distributional Result
Consider Table 2 and assume that all edges both in $G$ and $\hat{G}$ were placed uniformly at random. Then, given the number of true ($m_{\text{true}}$), estimated ($m_{\text{est}}$) and maximum total ($m_{\text{max}}$) edges, the number of true positives can be seen as the result of a simple random urn experiment with two colors of balls, say, blue and white: White balls correspond to adjacencies included in $G$, and blue balls are adjacencies not in $G$. A random causal discovery procedure will then metaphorically draw "balls" (i.e., edges) randomly without replacement, and some will be true positives (white), while others will be false positives (blue). Since the number of white balls ($m_{\text{true}}$), the number of draws ($m_{\text{est}}$) and the total number of balls ($m_{\text{max}}$) are all known a priori, the number of drawn white balls (true positive adjacencies) will by definition follow a hypergeometric distribution: $\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}}\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$.
Appendix B Computations for Table 3
@@
Below, we let $q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}$ be the $k$th quantile from $\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$. We also note, and use repeatedly below, that $\textit{TP}\operatorname{|}(m_{\text{max}},m_{\text{true}},m_{\text{est}})\sim\text{HyperGeom}(m_{\text{max}},m_{\text{true}},m_{\text{est}})$, and hence $\text{E}(\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})=\frac{m_{\text{est}}\cdot m_{\text{true}}}{m_{\text{max}}}$.
We will compute the expectation and quantiles for each of these four adjacency metrics: Recall, F1, negative predictive value (NPV), and specificity.
@@
We write recall as a function of $\textit{TP},m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
recall $=\frac{\textit{TP}}{\textit{TP}+\textit{FN}}$
$=\frac{\textit{TP}}{\textit{TP}+m_{\text{true}}-\textit{TP}}$
$=\frac{\textit{TP}}{m_{\text{true}}}.$
@@
$\text{E}(\text{recall}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{1}{m_{\text{true}}}\text{E}(\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$
@@
$\frac{1}{m_{\text{true}}}q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}.$
@@
We write the F1 score as a function of $\textit{TP},m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
F1 $=\frac{2\cdot\textit{TP}}{2\cdot\textit{TP}+\textit{FP}+\textit{FN}}$
$=\frac{2\cdot\textit{TP}}{2\cdot\textit{TP}+m_{\text{est}}-\textit{TP}+m_{\text{true}}-\textit{TP}}$
$=\frac{2\cdot\textit{TP}}{m_{\text{est}}+m_{\text{true}}}.$
@@
$\frac{2\cdot q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{est}}+m_{\text{true}}}.$
@@
We write the negative predictive value (NPV) as a function of $\textit{TP},m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
NPV $=\frac{\textit{TN}}{\textit{TN}+\textit{FN}}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\textit{TP}+\textit{FN}}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+TP}{m_{\text{max}}-m_{\text{est}}}.$
@@
$\text{E}(\text{NPV}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\text{E}(\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})}{m_{\text{max}}-m_{\text{est}}}$
@@
$\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{max}}-m_{\text{est}}}.$
@@
We write specificity as a function of $\textit{TP},m_{\text{max}},m_{\text{true}}$ and $m_{\text{est}}$:
specificity $=\frac{\textit{TN}}{\textit{TN}+\textit{FP}}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\textit{TP}}{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\textit{TP}+\textit{FP}}$
$=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\textit{TP}}{m_{\text{max}}-m_{\text{true}}}.$
@@
$\text{E}(\text{specificity}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})$ $=\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+\text{E}(\textit{TP}\operatorname{|}m_{\text{max}},m_{\text{true}},m_{\text{est}})}{m_{\text{max}}-m_{\text{true}}}$
@@
$\frac{m_{\text{max}}-m_{\text{est}}-m_{\text{true}}+q_{(k,m_{\text{max}},m_{\text{true}},m_{\text{est}})}}{m_{\text{max}}-m_{\text{true}}}.$
Appendix C Details About PC Algorithm Simulation Study
Let $G$ be the true graph and $\hat{G}$ be an estimate. We consider the following metrics:
Structural Hamming distance:
The structural Hamming distance (SHD) counts the number of edge reversals, removals and additions needed in order to transform $\hat{G}$ into $G$ [Tsamardinos et al., 2006].
Adjacency precision and recall:
Precision and recall computed from the adjacency/skeleton confusion matrix, as described in Section 2.
Orientation precision and recall:
Precision and recall computed from a conditional orientation confusion matrix. The conditional orientation confusion matrix is contructed as follows: For all edges that are both in $G$ and $\hat{G}$, each edge endpoint is classified as:
  • True positive if there is an arrowhead both in $G$ and $\hat{G}$.
  • True negative if there is a tail both in $G$ and $\hat{G}$.
  • False positive if there is an arrowhead in $\hat{G}$, but a tail in $G$.
  • False negative if there is a tail in $\hat{G}$, but an arrowhead in $G$.
Proportion recovered v-structures:
v-structures are node triples with the structure $A\to B\leftarrow C$ where $A$ and $C$ are non-adjacent. This metric counts how many such structures are correctly recovered by $\hat{G}$, divided by the total number of v-structures in $G$. If there are no v-structures in $G$, the value is set to 1 (interpreted as all structures being recovered).
Structural intervention distance:
The structural intervention distance (SID) counts the number of node pairs $(X_{i},X_{j})$ for which $\hat{G}$ does not provide a valid adjustment set for the total causal effect of $X_{i}$ on $X_{j}$ (assuming $G$ is the truth) [Peters and Bühlmann, 2015]. This property is only well-defined for DAG-DAG comparisons: If $\hat{G}$ is a CPDAG, it may have undirected edges and hence the total causal effects may not be identifed. In this case, one instead computes the SID for each DAG in the equivalence class specified by the CPDAG and reports the minimum and maximum values as bounds. For computing SIDs, we use the SID R package [Peters, 2023] with default settings [^4] Note that some cases result in warnings due to the estimated graph not being a proper CPDAG, or having large connected components. In both cases, SID is computed based on local expansions of the graphs..
We now provide a step-by-step example of how the negative controls are computed and used, following the template from Section 6. We focus on a single metric (SHD) and setting (dense). We proceed as follows:
1. Standard simulation study:
We draw 1000 random Erdős-Rényi type DAGs over 10 nodes, each with $m_{\text{true}}=30$ edges. From each DAG, we simulate 400 independent Gaussian observations with randomly drawn regression parameters and error variances [^5] We use default options for regression parameters and error variances from the simGausFromDAG() function in the causalDisco R package [Petersen, 2022].. For each dataset, we apply the PC algorithm using a test for vanishing partial correlations with significance level $\alpha=0.05$.
  1. (a)
    We compare each of the 1000 true DAGs with the corresponding estimated CPDAG provided by the PC algorithm by computing their SHDs. We find a mean SHD of $27.33$ with a 95% confidence interval of $(21;33)$ (based on the empirical quantiles).
  2. (b)
    We store the true DAGs as well as the distribution of the estimated number of edges across the 1000 applications of the PC algorithm. The number of estimated edges from PC range from 7 to 19 with a mean of $13.3$.
2. Negative control simulation:
We draw 1000 random CPDAGs over 10 nodes with number of edges independently sampled from the $m_{\text{est}}$ distribution from Step 1 (b).
3. Negative control evaluation:
For each of the 1000 negative controls, we compare with a true DAG from Step 1 by computing the SHD. We find a mean SHD of $30.23$ (95% CI: $26;36$).
4. Comparison:
We conduct pairwise comparisons of SHD values obtained by PC and negative controls. We find
$p=\frac{1}{1000}\sum_{i=1}^{1000}\mathbf{1}\left(\text{SHD}(G_{\text{true}}^{i},\hat{G}_{\text{PC}}^{i})\geq\text{SHD}(G_{\text{true}}^{i},\hat{G}_{\text{NC}}^{i})\right)=0.202$
and hence conclude that PC is not significantly different from the negative controls.
Appendix D Details About Sachs Data Application
For PC, LiNGAM, GES and BOSS, we apply the algorithms in Tetrad version 7.6.7 [Ramsey et al., 2018] with default settings. For NOTEARS, we report SHD and number of edges as provided in [Zheng et al., 2018].
The negative control DAGs are simulated as Erdős-Rényi type DAGs. The negative control CPDAGs are constructed by first simulating an Erdős-Rényi type DAG, and secondly constructing the CPDAG corresponding to its Markov equivalence class.
We use the 20 edge "truth" version of the Sachs dataset because that is also what was used to evaluate the NOTEARS algorithm in [Zheng et al., 2018]. We obtain both the ground truth graph and the Sachs dataset from this repository: https://github.com/cmu-phil/example-causal-datasets.